OpenStudy (lukecrayonz):

Shortcuts?:) Find dy/dx y=-3x^7

OpenStudy (anonymous):

that is simply the derivative... dy/dx= -21x^6

OpenStudy (lukecrayonz):

Yes I know. What is a shortcut? I just started learning it.

OpenStudy (lukecrayonz):

I know the formula, but is there a fast way to do it?

OpenStudy (anonymous):

A shortcut for monomials as such would be to use the power rule.

OpenStudy (anonymous):

y=ax^n dy/dx=anx^n-1

OpenStudy (lukecrayonz):

Okay that's what I figured a shortcut is, I just wanted to make sure it was an actual rule.

OpenStudy (lukecrayonz):

This only works for monomials?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

It works for any polynomial

OpenStudy (lukecrayonz):

I'm not on polynomial's yet but I'm sure I will be in a few minutes. How about this question? d/dx (18sqrt(x))

OpenStudy (anonymous):

\[18\sqrt{x}?\]

OpenStudy (anonymous):

\[18\sqrt(x)\] can also be rewritten as \[18x ^{1/2}\]